# Three meta-questions on infinite-domain Constraint Satisfaction Problems

### v1

@@ Appendix B $\omega$-categorical sandwiches for PCSPs
Let $\bar{s}_{1},\dots,\bar{s}_{m}$ be arbitrary tuples over $S$ of arity $k$ large enough so that, for every $\bar{s}\in S^{m}$, there exists $\ell\in[k]$ such that $\bar{s}$ equals the $\ell$-th entry in $(\bar{s}_{1},\dots,\bar{s}_{m})$ (viewed as a $k\times m$-matrix). Define the $m$-ary operations $\tilde{g_{1}}$ and $\tilde{g_{2}}$ by $\tilde{g_{1}}(x_{1},\dots,x_{m})\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=g_{1}(x_{i_{1}},\dots,x_{i_{n}})$ and $\tilde{g_{2}}(x_{1},\dots,x_{m})\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=g_{2}(x_{i_{n+1}},\dots,x_{i_{n+k}})$. Then the tuples $\tilde{g_{1}}(\bar{s}_{1},\dots,\bar{s}_{m})$ and $\tilde{g_{2}}(\bar{s}_{1},\dots,\bar{s}_{m})$ satisfy the same equalities and relations w.r.t. $\mathfrak{B}$ on all $d$-tuples. Since all relations of $\mathfrak{B}$ have arity $\leq d$ and $\mathfrak{B}$ is homogeneous, $\tilde{g_{1}}(\bar{s}_{1},\dots,\bar{s}_{m})$ and $\tilde{g_{2}}(\bar{s}_{1},\dots,\bar{s}_{m})$ lie in the same orbit. Thus, there are $\alpha_{1},\alpha_{2}\in\mathrm{Aut}(\mathfrak{B})$ such that
@@
Clearly, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$, we have ${\mathfrak{S}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. We show that there exists a finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ such that $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ is not finitely tractable. Since $\mathfrak{A}$ is $\omega$-categorical, Lemma B.2 guarantees the existence of some finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ with $|S|\geq 3$ such that $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ preserves $\neq$. Suppose that there is a finite sandwich $\mathfrak{D}$ of $\smash{({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})}$ that does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 3.1, there exists a cyclic operation $f$ in $\mathrm{Pol}(\mathfrak{D})$ of some arity $n$. Take any homomorphisms $g\colon{\mathfrak{S}}^{[d]}\rightarrow\mathfrak{D}$ and $h\colon\mathfrak{D}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Then, by setting $f^{\prime}(x_{1},\dots,x_{n})\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=h\circ f(g(x_{1}),\dots,g(x_{n})),$ we get a cyclic $n$-ary $f^{\prime}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}).$ By Lemma B.9 there is some $n$-ary $g\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that is pseudo-cyclic modulo $\mathrm{Aut}(\mathfrak{B})$. By Lemma B.1, this operation is cyclic. Take pairwise distinct elements $s_{1},s_{2},s_{3}\in S$. We now show that the cyclicity of $g$ contradicts the fact that it preserves $\neq$. To this end, we consider the following three cases:
$n\equiv 0\bmod 3$:

### v2

@@ Appendix B $\omega$-categorical cheeses for PCSPs
Let $\bar{s}_{1},\dots,\bar{s}_{m}$ be arbitrary tuples over $S$ of arity $k$ large enough so that, for every $\bar{s}\in S^{m}$, there exists $\ell\in[k]$ such that $\bar{s}$ equals the $\ell$-th entry in $(\bar{s}_{1},\dots,\bar{s}_{m})$ (viewed as a $k\times m$-matrix). Define the $m$-ary operations $\tilde{g_{1}}$ and $\tilde{g_{2}}$ by $\tilde{g_{1}}(x_{1},\dots,x_{m})\coloneqq g_{1}(x_{i_{1}},\dots,x_{i_{n}})$ and $\tilde{g_{2}}(x_{1},\dots,x_{m})\coloneqq g_{2}(x_{i_{n+1}},\dots,x_{i_{n+k}})$. Then the tuples $\tilde{g_{1}}(\bar{s}_{1},\dots,\bar{s}_{m})$ and $\tilde{g_{2}}(\bar{s}_{1},\dots,\bar{s}_{m})$ satisfy the same equalities and relations w.r.t. $\mathfrak{B}$ on all $d$-tuples. Since all relations of $\mathfrak{B}$ have arity $\leq d$ and $\mathfrak{B}$ is homogeneous, $\tilde{g_{1}}(\bar{s}_{1},\dots,\bar{s}_{m})$ and $\tilde{g_{2}}(\bar{s}_{1},\dots,\bar{s}_{m})$ lie in the same orbit. Thus, there are $\alpha_{1},\alpha_{2}\in\mathrm{Aut}(\mathfrak{B})$ such that
@@
Clearly, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$, we have ${\mathfrak{S}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Let $S\subseteq A$ with $|S|\geq 3$. Suppose that there is a finite cheese $\mathfrak{D}$ of $({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ that does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 3.1, there exists a cyclic operation $f$ in $\mathrm{Pol}(\mathfrak{D})$ of some arity $n$. Take any homomorphisms $g\colon{\mathfrak{S}}^{[d]}\rightarrow\mathfrak{D}$ and $h\colon\mathfrak{D}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Then, by setting $f^{\prime}(x_{1},\dots,x_{n})\coloneqq h\circ f(g(x_{1}),\dots,g(x_{n})),$ we get a cyclic $n$-ary $f^{\prime}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}).$ By Lemma B.7 there is some $n$-ary $g\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that is pseudo-cyclic modulo $\mathrm{Aut}(\mathfrak{B})$. By Lemma B.1, this operation is cyclic. Take pairwise distinct elements $s_{1},s_{2},s_{3}\in S$. We now show that the cyclicity of $g$ contradicts the fact that it preserves $\neq$. To this end, we consider the following three cases:
Case $n\equiv 0\bmod 3$: